Activity 1.1 Answer the guide questions to complete the table of domain and range of the following exponential functions, a.f(x) = 5* Is f(x)=5* defined at any values of x? What is the minimum value of f(x)? Can you determine the the maximum value of f(x)? Domain. Range Set Notation Interval Notation. b. f(x) = (-)" Is f(x) = ()** defined at any values of x? What is the minimum value of f(x)? Can you determine the maximum value of f(x)? Range Set Notation Interval Notation. Domain c. f(x) = (2)³ +3 Which is P(x)? Is a > 0 or a <0? Is P(x) linear? What is the value of h? Set Notation Interval Notation Domain Range d. g(x)=-(4)²-3+1 Which is P(x)? Is a > 0 or a <0? What is the value of h? Is P(x) linear? Set Notation Domain Range Interval Notation
1
Expert's answer
2021-10-29T01:43:03-0400

a) "f(x)=5^x"


Is "f(x)=5^x" defined at any values of "x" ? Yes.

What is the minimum value of "f(x)" ? Cannot be defined.

Can you determine the maximum value of "f(x)" No.



"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 & Domain & Range \\ \hline
 Set\ notation & \R & \{y\in \R|y>0\} \\
 \hdashline
 Interval\ notation & (-\infin,\infin) & (0,\infin)
\end{array}"

b) "f(x)=(\dfrac{1}{5})^{3x}"


Is "f(x)=(\dfrac{1}{5})^{3x}" defined at any values of "x" ? Yes.

What is the minimum value of "f(x)" ? Cannot be defined.

Can you determine the maximum value of "f(x)" No.



"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 & Domain & Range \\ \hline
 Set\ notation & \R & \{y\in \R|y>0\} \\
 \hdashline
 Interval\ notation & (-\infin,\infin) & (0,\infin)
\end{array}"

c) "f(x)=\dfrac{1}{3}(2)^{x}+3"


Which is "P(x)"? "x" Is "P(x)" linear? Yes.

Is "a>0" or "a<0" ? "a>0"

What is value of "h" ? "3"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 & Domain & Range \\ \hline
 Set\ notation & \R & \{y\in \R|y>3\} \\
 \hdashline
 Interval\ notation & (-\infin,\infin) & (3,\infin)
\end{array}"

d) "g(x)=-(4)^{2x-3}+1"


Which is "P(x)"? "2x-3" Is "P(x)" linear? Yes.

Is "a>0" or "a<0" ? "a<0"

What is value of "h" ? "1"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 & Domain & Range \\ \hline
 Set\ notation & \R & \{y\in \R|y<1\} \\
 \hdashline
 Interval\ notation & (-\infin,\infin) & (-\infin, 1)
\end{array}"

e) "h(x)=2(3)^{3x-1}+2"


Which is "P(x)"? "3x-1" Is "P(x)" linear? Yes.

Is "a>0" or "a<0" ? "a>0"

What is value of "h" ? "2"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 & Domain & Range \\ \hline
 Set\ notation & \R & \{y\in \R|y>2\} \\
 \hdashline
 Interval\ notation & (-\infin,\infin) & (2,\infin)
\end{array}"